# Loss-Oriented Ranking for Automated Visual Prompting in LVLMs

### v2

@@ 4.3.1 Comparison with Other Retrieval Methods
| AutoV | 38.7 | 52.1 | 32.9 | 41.3 |
@@
To analyze why ranking is more suitable for retrieving visual prompts, we compare it with three other retrieval methods. The experiments are conducted on LLaVA-7B. The retrieval pool is identical to that used in the main experiments.
Heuristic Fixed Retrieval. It involves a single fixed type of visual prompts, based on its predefined rules or heuristics. As shown in rows 3 to 6 of Table 2, we find that each visual prompt could potentially be the optimal solution for different benchmarks. For instance, achieves the best results on MMMU, while is optimal on MMVet, and there is a significant variation in the performance of different types of visual prompts across various benchmarks. In summary, heuristic retrieval offers simplicity in retrieval processes with the need for conducting benchmark-level multiple rounds of trials, consuming more resources.
Random Retrieval. It picks up one visual prompt from the candidate pool without any specific criteria or order. The th row of Table 2 lists its results. By comparing with the individual visual prompt, the performance of random retrieval is on par with or slightly superior across the three benchmarks. For instance, the random algorithm outperforms in MMMU by , in MMVet by . It demonstrates that by flexibly selecting from multiple visual prompt candidates, even random retrieval has the potential to outperform individual visual prompts, and the conclusion provides valuable insights for our approach.
MoE (or GateNet) Retrieval. It is a common technique for selecting the most relevant expert during model decision-making, where Gumbel Softmax enables differentiable weighting of the gating network [fedus2022switch]. In our scenario, the features of multiple visual prompts are input to GateNet to determine which one should be retained and fed into the LLM. We insert GateNet into the output of the projection layer and train it with the same amount of data and training time as AutoV. The results can be seen in the th row of Table 2. We observe a improvement with MoE compared to the baseline. However, the improvement compared to the random algorithm was only , lagging significantly behind AutoV by . We suspect that the Mixture-of-Experts architecture encounters a lack of generalization ability during the fine-tuning stage, which can easily lead to overfitting. Therefore, MoE retrieval may not be the optimal approach for visual prompt retrieval in LVLMs. It is important to adopt a more robust framework that integrates textual guidance to enhance overall performance effectively.
@@
Retrieval Visualization. We visualize the prompt distribution of retrieval strategies within - for significance. Figure 4 presents a bar graph comparing the retrieval decisions of GateNet and our AutoV on MMVet. The bar heights indicate the frequency of visual prompt retrieval, while the overlaid line plots depict the independent accuracy for each prompt. Our analysis reveals a notable discrepancy between the GateNet retrieval and the actual inference accuracy of visual prompts. For instance, GateNet retrieves , the second-most useful prompt, only times (38/218, of cases), despite its high accuracy. In contrast, AutoV exhibits stronger alignment with the standalone accuracy of visual prompts, and prefers to select and , the two prompts that own superior accuracy.
4.3.2 Scaling Law in the Visual Prompts Pool
| Pool Size | 0 | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 |
| Source | Baseline | +{API-L} | +{API-L} | +{API-L} | +{API-L} | +{RedCircle} | +{FGVP} | +{VTPrompt} | +{TVP} |
| MMMU | 36.3 | 37.3 | 37.5 | 38.3 | 38.3 | 38.6 | 38.7 | 38.7 | 38.9 |
| VizWiz | 50.0 | 50.3 | 50.8 | 51.0 | 51.9 | 52.1 | 52.1 | 52.6 | 52.6 |
| MMVet | 30.6 | 31.8 | 32.0 | 32.6 | 32.8 | 32.9 | 32.9 | 33.1 | 33.1 |
| Avg. | 39.0 | 39.8 | 40.1 | 40.7 | 41.0 | 41.2 | 41.3 | 41.5 | 41.6 |
Having more visual prompt candidates can offer a more diverse selection, enhancing the flexibility and adaptability of LVLMs. However, is it always better to have more visual prompts in AutoV? In the main experiment, we used the attention prompting images [api] from layers , , , and in the CLIP [CLIP] model, FGVP [yang2023fine], and RedCircle [shtedritski2023does] as inputs, with a total of visual prompts. Here, we consider extra scenarios ranging from prompts and conduct experiments on the MMMU, VizWiz, and MMVet. The results are listed in Table 3. First, we observe that when the visual prompts themselves can lead to performance improvement, the accuracy increases as their quantity grows, but the rate of accuracy improvement slows down as the number continues to rise. Taking the experiments on MMMU as an example, as the number of prompts increases, the performance gain remains at when increasing from to prompts. However, when further increasing to prompts, the improvement drops to . Then, even if a single visual prompt itself degrades accuracy, incorporating it into the candidate pool does not hurt overall performance and can even lead to slight improvements, thanks to the inherent retrieval capability of AutoV. Taking the results on VizWiz as an example: although both RedCircle and FGVP individually show negative gains ( and , respectively), including them in the candidate pool leads to an overall improvement. In summary, AutoV exhibits notable robustness to variations in visual prompt quality, consistently preserving or enhancing overall performance across diverse settings. We recommend selecting known high-quality visual prompts as candidates, while limiting to fewer than instances to strike the balance between performance and efficiency.
4.3.3 The Transferability of AutoV
@@
4.3.4 Statistical Significance Analysis
To rigorously evaluate the reliability of our reported improvements, we conduct significance testing on the results in Table 1 (LLaVA-1.5 7B) across multiple random seeds. Taking the MMMU benchmark as an example, a paired t-test [hsu2014paired] confirms that improvement of AutoV over the baseline is highly significant (mean , ), with an extremely small -value, indicating that the observed gain is unlikely to arise from random chance. This supports our claim that AutoV consistently enhances VLM performance across repeated runs. Detailed statistical tables and analyses are provided in the Appendix E.
4.3.5 The Efficiency of AutoV
While we enhance LVLM performance through AutoV, we uphold efficiency optimizations to ensure practicality. Initially, training a high-quality AutoV on the LLaVA-7B model within epochs only requires hours utilizing NVIDIA A-GB GPUs. Furthermore, because of its transferability, training additional AutoV instances for each vision-language model is unnecessary, and we can generate the training samples via smaller models. Lastly, although AutoV introduces additional TFLOPs during inference (e.g., multiple image pre-processing and visual feature extraction), we can reduce latency by downsizing the final similarity matrix and single-prompt selection before LLM processing. As a result, the inference latency of the LLaVA-7B model increases by just ms (from ms to ms), while yielding a gain in average accuracy.
4.3.6 Qualitative Visualization
To validate the effectiveness of our retrieval strategy, we visualize an example sourced from LLaVA-Bench within {API-L-Dark, RedCircle, API-L-White, FGVP}. As shown in Figure 5, we list the output results of AutoV when we enter different queries. When faced with a question that focuses on the global context information, AutoV chooses (API-L-Dark), where the black mask removes the background. While referring to the ”intended effect” in the image, AutoV selects (RedCircle), which directly uses red circles to highlight the weird spots. More visualization retrieval cases can be found in the Appendix F.
@@ 5 Conclusion
In this paper, we introduce AutoV, an adaptive visual prompting framework that retrieves optimal visual prompts tailored to textual queries and image contexts for large vision-language models. By leveraging a lightweight ranking network trained with reward-based supervision, AutoV consistently outperforms existing heuristic prompting methods. Extensive experiments across multiple LVLM architectures and benchmarks demonstrate its effectiveness and generality. Overall, AutoV provides a practical and robust solution for enhancing LVLMs’ performance through adaptive visual prompting. Future work explores integrating AutoV with more advanced prompt generation techniques or extending its framework to broader multimodal settings beyond vision-language tasks.
@@
Contents
A Details of Training Data
| Method | CVBench | BLINK | RealWorldQA | MMVP | COCOCap | RefCOCO+ | ImageNet-1K | Avg. |
| LLaVA-OneVision (7B) | ||||||||
| Base Model | 79.6 | 48.2 | 66.0 | 60.7 | 61.1 | 68.8 | 92.3 | |
| API (ECCV24) | 80.7 | 49.4 | 66.4 | 61.5 | 61.9 | 69.4 | 92.6 | |
| AutoV | 83.0 | 51.6 | 68.9 | 64.2 | 63.7 | 71.8 | 95.0 | |
| Qwen2.5-VL (7B) | ||||||||
| Base Model | 80.6 | 56.4 | 68.5 | 61.3 | 40.7 | 72.6 | 94.7 | |
| API (ECCV24) | 81.1 | 57.3 | 69.2 | 61.7 | 41.4 | 73.5 | 95.4 | |
| AutoV | 83.0 | 59.0 | 71.1 | 64.3 | 43.3 | 75.2 | 97.1 | |
@@ A.2 The Composition of Training Data
B Evaluation Benchmarks
We conduct comprehensive experiments on fifteen widely used vision-language benchmarks.
@@ B Evaluation Benchmarks
C Implementation Details
(1) Data generation stage: LLaVA-v1.5 is consistently adopted as the reference LVLM for generating loss due to resource limitations. (2) AutoV training stage: The LLaVA series is trained using DeepSpeed [rasley2020deepspeed] ZeRO2 with A100-40G GPUs within epochs. The batch size (with accumulation) is set to and the learning rate is -. For convenience, Qwen2.5-VL, InternVL2, Gemini-1.5-Pro, GPT-4o, and Qwen-VL-Max directly adopt the retrieval strategy in AutoV pre-trained from LLaVA-OneVision.
D More Main Experiments
In this section, we present the extended results that complement the main experiments reported in Section 4. Specifically, we provide detailed quantitative results and additional analyses of AutoV on a broader range of benchmarks, including CVBench, BLINK, Real-World QA, and MMVP, as well as classical tasks such as COCO2017 Captioning, RefCOCO+ Grounding, and ImageNet-1K Classification. All experiments are conducted using LLaVA-OneVision 7B and Qwen2.5-VL 7B as base models. The results consistently demonstrate that AutoV effectively enhances multimodal understanding and reasoning across diverse domains, confirming its model-agnostic generalization capability.
@@ D More Main Experiments
E Statistical Significance Analysis
@@
F More Visualization Results
F.1 Cases of Training Data
@@
F.2 Retrieval Strategy of AutoV
To further demonstrate the effectiveness of our retrieval strategy, Figure 8 provides additional visualizations of ten representative cases beyond those presented in the main text, showing how AutoV retrieves the most suitable visual prompt according to different textual queries. These examples are sourced from different benchmarks, covering multiple visual prompt sets, including {API-L-Dark, RedCircle, API-L-White, FGVP}. Best viewed in color.

### v3

@@ 5.1 Comparison with Other Retrieval Methods
| AutoV List-wise ranking | 38.0 | 51.3 | 32.4 | 40.6 |
| AutoV Pair-wise ranking | 38.7 | 52.1 | 32.9 | 41.3 |
@@
To analyze why pair-wise ranking is more suitable for visual prompt retrieval, we compare it with several representative retrieval strategies. All experiments are conducted on LLaVA-7B with the retrieval pool utilized in the main experiments.
1. Heuristic Fixed Retrieval. This strategy selects a single predefined visual prompt based on manually designed rules. As shown in Rows 3–6 of Table 3, different prompts exhibit distinct performance profiles across benchmarks. For instance, performs best on MMMU, while is optimal on MMVet, indicating that no single prompt engineering is universally effective.
2. Random Retrieval. Randomly selecting yields a +0.7% average gain over the baseline and performs on par with or slightly better than individual fixed prompts. This suggests that prompt retrieval itself provides benefits.
3. Regression Retrieval. This strategy directly predicts the absolute loss score of each visual prompt and selects the lowest one. As shown in Table 3, regression improves the baseline by , but performs worse than both MoE and ranking-based methods. This suggests that learning absolute loss is challenging due to small performance gaps among prompts and inherent noise in supervision signals, making it difficult to reliably distinguish the optimal prompt.
4. MoE (GateNet) Retrieval. We further compare with a Mixture-of-Experts (MoE) style retrieval mechanism, where a lightweight gating network (implemented with Gumbel Softmax [fedus2022switch]) learns to select among multiple visual prompts in a differentiable manner. GateNet is inserted after the projection layer and trained under the same data and training budget as AutoV. As shown in Row 9 of Table 3, MoE improves the baseline by . While this demonstrates the benefit of learnable selection, its gain over random retrieval is modest (+0.4%), and it remains behind AutoV (pair-wise ranking).
5. List-wise Ranking. We further adopt a list-wise ranking objective that models the relative ordering of all candidate prompts simultaneously. List-wise ranking achieves consistent gains (+1.6%), demonstrating the benefit of structured preference learning over regression or soft gating. However, compared to pair-wise ranking, its improvement is more limited. We conjecture that list-wise optimization requires estimating a full permutation over candidates, which introduces higher optimization complexity and sensitivity to noisy comparisons.
6. AutoV. Overall, these comparisons suggest that effective visual prompt retrieval requires (i) query-aware adaptability beyond random diversity, (ii) explicit relative preference modeling beyond soft gating, and (iii) structured supervision beyond list-wise ranking. Pair-wise ranking naturally satisfies these properties, which explains its consistent superiority in Table 3.
5.2 Retrieval Visualization
@@
We visualize the prompt distribution of retrieval strategies within - for significance. Figure 4 presents a bar graph comparing the retrieval decisions of GateNet and our AutoV on MMVet. The bar heights indicate the frequency of visual prompt retrieval, while the overlaid line plots depict the independent accuracy for each prompt. Our analysis reveals a notable discrepancy between the GateNet retrieval and the actual inference accuracy of visual prompts. For instance, GateNet retrieves , the second-most useful prompt, only times (38/218, of cases), despite its high accuracy. In contrast, AutoV exhibits stronger alignment with the standalone accuracy of visual prompts, and prefers to select and , the two prompts that own superior accuracy.
5.3 The Scaling Law of AutoV
| Pool Size | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 |
| Source | API-L15 | API-L20 | API-L22 | API-L23 | RedCircle | FGVP | VTP | TVP |
| MMMU | 37.3 | 37.5 | 38.3 | 38.3 | 38.6 | 38.7 | 38.7 | 38.9 |
| VizWiz | 50.3 | 50.8 | 51.0 | 51.9 | 52.1 | 52.1 | 52.6 | 52.6 |
| MMVet | 31.8 | 32.0 | 32.6 | 32.8 | 32.9 | 32.9 | 33.1 | 33.1 |
| Avg. | 39.8 | 40.1 | 40.7 | 41.0 | 41.2 | 41.3 | 41.5 | 41.6 |
Having more visual prompt candidates increases selection diversity and potentially improves the adaptability of LVLMs. However, is more always better? Here, we extend the candidate pool from to prompts (Table 4). (1) We observe a clear diminishing-return effect. When prompts are individually beneficial, increasing their number initially improves accuracy, but the marginal gain decreases as the pool grows. For example, on MMMU, the gain remains when increasing from to prompts, but drops to when expanding from to prompts. (2) AutoV is largely agnostic to the quality of individual visual prompts, and its effectiveness does not depend on carefully curated prompt candidates. On VizWiz, although RedCircle and FGVP yield negative gains ( and ), incorporating them still leads to a improvement, demonstrating its robustness. (3) Moreover, even when encountering unseen prompt types during training, AutoV can effectively index suitable candidates, yielding an additional average improvement on the last two visual prompts.
Overall, AutoV remains stable under varying prompt quality and pool sizes, consistently preserving or improving performance. In practice, we recommend using high-quality prompts while limiting the pool to fewer than candidates to balance effectiveness and efficiency.
5.4 Study on the LLM Layer Selection
| Method | MMMU | VizWiz | MMVet |
| LLaVA-1.5-7B | 36.3 | 50.0 | 30.6 |
| Part1: LLM layer selection | |||
| 24th layer | 38.1 | 51.6 | 32.3 |
| 12th layer | 38.4 | 51.7 | 32.3 |
| 0th layer (Default) | 38.7 | 52.1 | 32.9 |
| Part2: Pre-filtering strategy | |||
| When 6 Candidates | |||
| w/o pre-filtering | 38.3 | 52.0 | 32.6 |
| w/ pre-filtering (Default) | 38.7 | 52.1 | 32.9 |
| When 8 Candidates | |||
| w/o pre-filtering | 38.4 | 52.2 | 32.7 |
| w/ pre-filtering (Default) | 38.9 | 52.6 | 33.1 |
Findings from FastV [chen2024image] and LLaVA-Mini [zhang2025llava] validate that cross-modal interactions mainly occur in shallow layers. This enables the textual input to better guide visual selection. Part 1 of Table 5 shows that the 0th layer outperforms intermediate or bottom layers, validating our design choice.
5.5 Study on the Pre-filtering Step
Part 2 of Table 5 shows consistent improvements, especially with more prompts. High-recall pre-filtering removes only the most dissimilar candidates, with a negligible risk of discarding optimal prompts (0.5% in our statistics), while substantially reducing the presence of harmful prompts.
5.6 The Transferability of AutoV
@@
5.7 Statistical Significance Analysis
To evaluate the reliability of our improvements, we perform paired t-tests [hsu2014paired] across multiple random seeds for all benchmarks in Table 1 (LLaVA-1.5 7B). Taking MMMU as an example, AutoV achieves a mean improvement of over the baseline with , indicating statistical significance rather than random fluctuation. Similar significance is consistently observed across other benchmarks. Detailed statistical results are provided in Appendix F.
5.8 The Efficiency of AutoV
| Factor | Base Setting | Expanded Setting | ||
| Extra (T) | Total (T) | Extra (T) | Total (T) | |
| Pool Size | 0.74 (4) | 5.08 | 1.62 (8 ) | 5.96 |
| Model Size | 0.74 (7B) | 5.08 | 0.74 (13B ) | 8.77 |
| Vision Token | 0.74 (576) | 5.08 | 1.48 (1152 ) | 6.59 |
AutoV improves LVLM performance while remaining lightweight and practical. Training AutoV on LLaVA-7B takes only 6 hours on 8 A100GB GPUs (40 epochs). Thanks to its transferability, no retraining is required for new LVLMs, and training data can be generated with smaller models. At inference, AutoV encodes additional images and ranks candidates (), without extra decoder cost. For LLaVA-7B (576 visual, 20 text tokens), the encoder, ranking-net, and decoder require 0.16T, 0.09T, and 4.30T FLOPs, respectively: LLM decoding clearly dominates. Under the base setting (), AutoV adds only 0.74T FLOPs (Table 7). Larger pools or higher resolution increase vision-side cost nearly linearly, yet remain negligible relative to decoder computation. The overhead is independent of model scale (7B vs. 13B). In practice, latency increases by just 6.9 ms (254.8 ms 261.7 ms) due to compact similarity computation and early prompt selection, yielding a +2.4% average accuracy gain.
5.9 Qualitative Visualization
To validate the effectiveness of our retrieval strategy, we visualize an example sourced from LLaVA-Bench within {API-L-Dark, RedCircle, API-L-White, FGVP}. As shown in Figure 5, we list the output results of AutoV when we enter different queries. When faced with a question that focuses on the global context information, AutoV chooses (API-L-Dark), where the black mask removes the background. While referring to the "intended effect" in the image, AutoV selects (RedCircle), which directly uses red circles to highlight the weird spots. More visualization retrieval cases can be found in the Appendix G.
@@ 6 Conclusion
In this paper, we introduce AutoV, an adaptive framework that retrieves optimal visual prompts for textual queries in LVLMs. Leveraging a lightweight ranking network trained with reward-based supervision, AutoV consistently outperforms prompt engineering methods. Notably, its effectiveness does not rely on the intrinsic quality of individual visual prompts, nor is it constrained by their standalone performance. Extensive experiments on multiple LVLMs and benchmarks demonstrate its effectiveness, robustness, and generality.
@@
Appendix A Details of Training Data
@@ A.2 The Composition of Training Data
Appendix B Evaluation Benchmarks
We conduct comprehensive experiments on fifteen vision-language benchmarks.
@@ Appendix B Evaluation Benchmarks
Appendix C Implementation Details
(1) Data generation stage: LLaVA-v1.5 7B is consistently adopted as the reference LVLM for generating loss due to resource savings. In total, 40K samples require loss annotation. With standard single A100-40G GPU inference, the entire process can be completed within half a day, making the data generation stage computationally manageable.
(2) AutoV training stage: The LLaVA series is trained using DeepSpeed [rasley2020deepspeed] ZeRO2 with A100-40G GPUs within epochs. The batch size (with accumulation) is set to and the learning rate is -. For convenience, Qwen2.5-VL, InternVL2, Gemini-1.5-Pro, GPT-4o, and Qwen-VL-Max directly adopt the retrieval strategy in AutoV pre-trained from LLaVA-OneVision.
Appendix D Reliability of Language-modeling Loss
We mitigate language priors using a low-variance criterion, where small loss variation across images indicates weak visual dependence. For instance, we conduct a query-only evaluation on the training set, where accuracy drops from before filtering to after filtering, indirectly indicating stronger visual grounding in the filtered data. Besides, we adopt the GPT-4o to judge language-prior cases, where the proportion decreases from to after filtering. Therefore, these results provide stronger empirical support.
| Models | VizWiz | LLaVA | Models | VizWiz | LLaVA |
52.8 | 68.0 |
64.6 | 85.8 |
| Policy from LLaVA-OneVision (0.5B) | |||||
| + AutoV | 56.9 | 69.7 | + AutoV | 68.9 | 87.0 |
| Policy from LLaVA-OneVision-1.5 (4B) | |||||
| + AutoV | 61.4 | 70.3 | + AutoV | 71.1 | 88.1 |
Appendix E Analysis of the Teacher (Strategy) Model Scale
We study the effect of teacher scale using smaller LLaVA-OneVision series models (0.5B and 4B). As shown in Table 8, AutoV consistently improves performance on GPT-4o, e.g., on VizWiz and on LLaVA, even with the smallest B teacher. Therefore, AutoV is robust to teacher scale, while larger teachers provide slightly higher gains, enabling flexible efficiency–performance trade-offs for users.
@@ Appendix E Analysis of the Teacher (Strategy) Model Scale
Appendix F Statistical Significance Analysis
@@
Appendix G More Visualization Results
G.1 Cases of Training Data